There are 3 consecutive integers. Three times the first is the same as 17 more Than 2 times the third. What is the second integer
Answer by checkley77(12844)

x, x+1 & x+2 are the three integers.
3x=2(x+2)+17
3x=2x+4+17
3x-2x=21
x=21 for the smallest integer.
21+1=22 for the middle integer.
Proof:
3*21=2(21+2)+17
63=2*23+17
63=46+17
63=63